# Accelerated Ostwald ripening by chemical activity

### v1

@@ III.3 Quasistatic growth
Compared to the passive scenario of Sec. II.3 with only sticky particles, the saturation of the sticky species has decreased from $\varphi^{\infty}$ to $\varphi_{\mathrm{s}}^{\infty}$. Does this loss of saturation still allow for accelerated Ostwald ripening?
III.4 Droplet dynamics
In this section, we find the quasistatic volume-fraction fields, and obtain the relevant expressions for the focal-droplet radius dynamics under the effect of reactions.
We find the solution to Eqs. (15),
$\phi_{\mathrm{s}\mathrm{g}}(r;R)+\phi_{\mathrm{ns}\mathrm{g}}(r;R)$
$\quad=\varphi^{\infty}-\left[\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)\right]\frac{R}{r},$ (17)
$k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}(r;R)-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}(r;R)$
$\quad=\left[k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)\right]\frac{R}{r}e^{-(r-R)/\ell}.$ (18)
Eq. (17) is reminiscent of Eq. (7), as the total volume fraction of proteins is not affected by the interconversion reaction, and hence its profile is simply diffusive as in the passive case. In Eq. (18), we identified explicitly the reactive lengthscale — the distance that a particle diffuses in between reactions [3],
$\ell=\sqrt{\frac{D_{\mathrm{g}}}{k_{\mathrm{ns}\to\mathrm{s}}+k_{\mathrm{s}\to\mathrm{ns}}}}.$ (19)
It sets the distance over which the reactions equilibrate ($s(\phi_{\mathrm{s}\mathrm{g}}(r;R),\phi_{\mathrm{ns}\mathrm{g}}(r;R))=0$). As mentioned above, for the assumption of fast surface thermodynamics to hold, we need $\ell\gg\kappa_{\mathrm{s}}^{1/2}$. In Sec. IV we will investigate two regimes, $R\ll\ell$ and $R\gg\ell$. Upon isolating the sticky and nonsticky volume-fraction profiles from the linear combinations given in Eq. (17) and Eq. (18), we find the gas-phase sticky flux into the droplet,
$J_{\mathrm{s}\mathrm{g}}(R)$ $=$ $-\frac{D_{\mathrm{g}}}{R}[\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)]$ (20)
$-\frac{D_{\mathrm{g}}}{\ell}\frac{k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)-k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}.$
An identical expression for $J_{\mathrm{ns}\mathrm{g}}(R)$ is obtained upon replacing $\mathrm{s}\leftrightarrow\mathrm{ns}$. For growing droplets, the ratio $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}/\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ between sticky and nonsticky volume fractions in the gas phase just outside the droplet, is smaller than the one dictated by the reactions, $k_{\mathrm{ns}\to\mathrm{s}}/k_{\mathrm{s}\to\mathrm{ns}}=\varphi_{\mathrm{s}}^{\infty}/\varphi_{\mathrm{ns}}^{\infty}$. Thus, the second term in Eq. (20), resulting from the reactions, joins the first term arising from passive Ostwald ripening (Eq. (8)) in bringing material into the droplet, and vice versa for shrinking droplets. This fact is illustrated in Fig. 2, and it is the possible source of accelerated Ostwald ripening.
Figure 2: Illustrative concentration profiles between a small and a large droplet in the passive and active cases, where $\Delta\varphi_{\mathrm{s}}=\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)$ is the excess saturation of sticky particles in the dilute phase. In the passive case, close to each of the droplets, the concentration profile is hyperbolic (dashed black curve), bringing material to the large droplet and away from the small one, depending on the dilute phase concentration between the droplets. In the active case, reactions create nonsticky material, thus overall lowering the sticky volume fraction in the dilute phase. However, the reactions may make the sticky profile steeper in the vicinity of droplets, thus increasing the flux of material from the small to the large droplet. In both cases, the indicated proportionalities for $\Delta\varphi_{\mathrm{s}}$ hold for both the smaller and larger droplet, up to a change of sign for the large droplet. The full expressions for the volume fraction profiles can be found in the SI Appendix, Sec. II.
As mentioned above, linearizing Eq. (13) in the dense phase is much more intricate, and we consign these calculations to the SI Appendix, Sec. II. In contrast to the dilute phase, the dense phase is almost entirely composed of sticky proteins. This means that reactions inside the droplet hinder its growth, as their net effect is to convert sticky material into the nonsticky form, $s_{\mathrm{l}}\equiv s(\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}})>0$. (We first mention that a phenomenon akin to Turing instability [40] might occur for certain diffusive coupling values between sticky and nonsticky species. In this case, droplets may not be stable once $R(t)$ is larger than the characteristic length for the patterns. We do not explore this phenomenon further, and we assume that the parameter values are such that it is avoided, and hence the volume fraction profiles inside the droplets are stable.) For small droplets, the amount of converted material per time is proportional to droplet volume. However, once the droplet is larger than the distance a particle typically diffuses between reactions inside the droplet, the reactions equilibrate at the core of the droplet, [^4] We assume that the equilibrated volume fraction of sticky species, given by $s(\phi_{\mathrm{s}\mathrm{l}}(0;R),\phi_{\mathrm{ns}\mathrm{l}}(0;R))=0$ and $\phi_{\mathrm{s}\mathrm{l}}(0;R)+\phi_{\mathrm{ns}\mathrm{l}}(0;R)=\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}(R)+\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}(R)$ is still in between the binodal and spinodal lines of the free energy, so bubbles do not form [9]; see SI Appendix, Sec. II. so the net loss of sticky material occurs only at its interface and hence the total conversion rate is proportional to droplet surface area. By contrast, Ostwald ripening, which opposes this loss of material, occurs with a rate proportional to droplet radius, whereas the gas-phase reactions contribute additional ripening at a rate proportional to the area as well. These competitions produce a set of nonequilibrium behaviors which are explored below and summarized in Fig. 3.
We conclude this section by calculating the droplet growth rate. By equating Eq. (5) and Eq. (12), we find to leading order in small gas-phase volume fractions,
$\frac{k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)-k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}-\frac{s_{\mathrm{l}}g(R/\ell)}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}$
$=\frac{\ell}{R}[\varphi^{\infty}_{\mathrm{ns}}-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)]$ , (21)
from which we can infer $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ [37]. Details of this calculations, as well as the functional form of $g(x)$ (originating from the dense-phase calculations), can be found in SI Appendix, Sec. II. On the other hand, expressing Eq. (5) and Eq. (12) for the total volume fraction field $\phi_{\mathrm{s}}+\phi_{\mathrm{ns}}$, we find
@@ III.4 Droplet dynamics
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R}\left[\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)\right].$ (22)
@@
Eq. (22) is a key result — upon comparison with Eq. (9), it emphasizes that regardless of the activity, this system is bound by the scaling laws of conventional Ostwald ripening. As mentioned previously, this is guaranteed by the conservation law of $\phi_{\mathrm{s}}+\phi_{\mathrm{ns}}$, a volume fraction field which is not directly affected by activity. Therefore, in our search for acceleration, we understand that the scaling laws expected from Ostwald ripening are the best case scenario, requiring small $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, and our only hope is to increase the rate coefficient. Otherwise, in the “worst case scenario” of large $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, Ostwald ripening might be completely prevented. The alternative possible outcomes we capture below include size control [3] and arrested ripening (that is, no shrinkage of small droplets) [6].
With this we have laid out our theory, and are prepared to list our results. In Sec. II.4, only the lengthscale $\lambda$ associated with surface tension was relevant, and was assumed to be much smaller than the typical droplet size $R$ during ripening. Here, $\ell$ of Eq. (19) is another, larger lengthscale to which $R$ must be compared. Thus, we split our discussion into small ($R\ll\ell$) and large ($R\gg\ell$) droplets. Detailed calculations can be found in SI Appendix, Sec. II.
IV Results
With the theoretical setup of Sec. III, we proceed to obtain explicitly the dynamics of a focal droplet. The reactions complicate the standard Ostwald ripening calculations of Sec. II [24]. First, the distance a particle diffuses before reacting splits droplet growth into two regimes, which we study in separate sections. Furthermore, the reactions within the liquid hinder Ostwald ripening, requiring special attention to identify the parameter values for which growth is permitted. Recent reviews on the known possible outcomes can be found in Refs. [41, 1].
IV.1 Small droplet
For small droplets, we compute the undetermined $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ using Eq. (21) to leading order in small dilute-phase density $\varphi^{\infty}\ll 1$ and radius $R/\ell\ll 1$, and keep the leading order term containing the net reaction rate inside the droplet, $s_{\mathrm{l}}$,
$\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)$ $=$ $\varphi^{\infty}_{\mathrm{ns}}-\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}[\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)]\frac{R}{\ell}$ (23)
$+\frac{1}{3}\frac{s_{\mathrm{l}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}\left(\frac{R}{\ell}\right)^{2},$
where $g(x)=x/3$ in this limit. Equipped with $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)$, Eq. (22) becomes

### v2

@@ Abstract
To enable phase separation, the system is prepared in some initial, slightly supersaturated concentration denoted $\varphi^{\infty}\gtrsim\phi_{\mathrm{s}\mathrm{g}}^{\circ}$. This so-far passive solution is known to undergo Ostwald ripening; cf. Fig. 1(a). As the saturation decreases during droplet growth, since a higher dilute-phase concentration $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)$ surrounds smaller droplets due to Laplace pressure, a concentration gradient diffusively transports material from small to large droplets. This is summarized by the radius growth rate equation for a focal droplet [24],
@@
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R}\left[(\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ})-\frac{\lambda_{\mathrm{s}\mathrm{g}}}{R}\right],$ (1)
where $D_{\mathrm{g}}$ is the dilute-phase diffusion constant. Upon inspecting the size distribution of thus growing and shrinking droplets, the Lifshitz-Slyozov scaling law emerges[24],
$\langle R(t)\rangle^{3}-\langle R(0)\rangle^{3}=\frac{4}{9}D_{\mathrm{g}}\lambda_{\mathrm{s}\mathrm{g}}t.$ (2)
Are active sticky-to-nonsticky interconversion reactions capable of accelerating this growth law?
To a sticky-only solution of (super-) saturation $\varphi^{\infty}$, we add a catalyst (e.g., a kinase) that converts sticky proteins into nonsticky ones, and vice versa by another catalyst (e.g., a phosphatase); cf. Fig. 1(b). For example, in an aqueous solution, the sticky and nonsticky species might be interconvertible hydrophobic and hydrophilic compounds. Since these reactions are independent of the underlying thermodynamics of phase separation, they normally consume fuel whose concentration we assume is fixed [1, 15]. We define the total reaction rate, $s(\phi_{\mathrm{s}},\phi_{\mathrm{ns}})$, as a function of the local particle concentrations, with the convention that $s>0$ implies net conversion of sticky molecules to nonsticky. For example, a purely first-order reaction mechanism takes the form $s(\phi_{\mathrm{s}},\phi_{\mathrm{ns}})=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}}$, where $k_{i\to j}$ are the conversion rates, with associated stoichiometric coefficients $\upsilon_{\mathrm{s}}=-1$ and $\upsilon_{\mathrm{ns}}=1$.
By nonsticky we mean that these particles are inert, meaning that phase separation is still driven by the remaining sticky species. We denote the coexistence concentrations of the dense phase (sticky-rich, “liquid”) by $\{\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\}$ and of the dilute phase (sticky-poor, “gas”) by $\{\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}\}$ around droplets. When the two phases in contact are in equilibrium, we now require equal sticky chemical potentials, nonsticky chemical potentials, and pressures.
In accordance with the Gibbs phase rule [27], these three conditions are insufficient to determine all four concentrations. Therefore, we next impose material conservation. During coarsening, the net amount of sticky material joining the droplet must track with the net amount of nonsticky material being expelled [28]. Explicitly,
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{J_{\mathrm{s}\mathrm{l}}(R)-J_{\mathrm{s}\mathrm{g}}(R)}{\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}(R)-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)}=\frac{J_{\mathrm{ns}\mathrm{l}}(R)-J_{\mathrm{ns}\mathrm{g}}(R)}{\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}(R)-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)},$ (3)
[…]
The fluxes can be computed using rigorous approximations. Specifically, within linear-irreversible thermodynamics [29], the concentrations respond linearly to gradients in the chemical potentials, which are determined by the concentration profiles. These concentrations, in turn, are affected by the active interconversion reactions occurring in the solution. We summarize below the derivation and its results, and defer all rigorous calculations to Sec. SII. In what follows, we assume that droplet growth is quasistatic and that the activity does not modify surface thermodynamics [30, 31].
In the dilute phase, we suppose that the free energy assumes the form of an ideal gas, so Fickian diffusion applies. Likewise, we approximate the reactions as effectively first-order, $s(\phi_{\mathrm{s}\mathrm{g}},\phi_{\mathrm{ns}\mathrm{g}})\simeq k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}$, as only the lowest-order reaction mechanisms take place with a nonzero rate in the small concentration limit. (We generalize this in Sec. SII D.) The effective rate constants $k_{\mathrm{s}\to\mathrm{ns}}$ and $k_{\mathrm{ns}\to\mathrm{s}}$ can be derived from the full reaction mechanism $s(\phi_{\mathrm{s}},\phi_{\mathrm{ns}})$, and they depend on fuel and catalyst concentrations. Due to the presence of reactions, we note that the total superaturation $\varphi^{\infty}$ is now redistributed among sticky and nonsticky species according to chemical equilibrium,
$\varphi_{\mathrm{s}}^{\infty}=\frac{k_{\mathrm{ns}\to\mathrm{s}}\varphi^{\infty}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}\text{\,\,\,and\,\,\,}\varphi_{\mathrm{ns}}^{\infty}=\frac{k_{\mathrm{s}\to\mathrm{ns}}\varphi^{\infty}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}.$ (4)
Does this loss of sticky-particle saturation, $\varphi_{\mathrm{s}}^{\infty}<\varphi^{\infty}$, still allow for accelerated Ostwald ripening?
The competition between diffusion and reactions in the dilute phase introduces a lengthscale reflecting the diffusive distance over which reactions occur,
$\ell_{\mathrm{g}}=\sqrt{\frac{D_{\mathrm{g}}}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}}}.$ (5)
With its aid, upon finding the dilute-phase steady profiles from the reaction-diffusion equation, the material fluxes read
$J_{\mathrm{s}\mathrm{g}}(R)=$ $-\frac{D_{\mathrm{g}}}{R}[\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)]$
$-\frac{D_{\mathrm{g}}}{\ell_{\mathrm{g}}}\frac{k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)-k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}},$ (6)
where $J_{\mathrm{ns}\mathrm{g}}(R)$ is also given by Eq. (6) upon replacing $\mathrm{s}\leftrightarrow\mathrm{ns}$. The first term on the right side of Eq. (6) represents passive diffusion due to the concentration gradient from $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}$ at the interface to $\varphi_{\mathrm{s}}^{\infty}$ in the bulk. The second term due to active reactions steepens the concentration gradient around the interface for both growing and shrinking droplets. Therefore, the effect of reactions is to further promote ripening. This fact is illustrated in Fig. 2 and serves as the basis for the acceleration we report.
Figure 2: Illustrative concentration profiles between a small and a large droplet in the passive and active cases, where $\Delta\varphi_{\mathrm{s}}(r)=\phi_{\mathrm{s}}(r)-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}$ is the excess saturation of sticky particles in the dilute phase. In the passive case, close to each of the droplets, the concentration profile is hyperbolic (dashed black curve), bringing material away from the small droplet and to the large one. In the active case, reactions create nonsticky material, thus overall lowering the sticky concentration in the dilute bulk. However, the reactions make the sticky profile steeper in the vicinity of droplets, thus increasing the flux of material from the small to the large droplet. In both cases, the indicated proportionalities for $\Delta\varphi_{\mathrm{s}}$ hold for both the smaller and larger droplet, up to a change of sign for the large droplet. The full expressions for the concentration profiles can be found in Sec. SII F.
The dense phase is almost entirely composed of sticky proteins. This means that reactions inside the droplet can only hinder its growth, as their net effect is to convert sticky material into the nonsticky form, $s_{\mathrm{l}}\equiv s(\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}})>0$. When the droplets are still small, the total amount of converted material per unit time is proportional to droplet volume. However, once the droplet is larger than $\ell_{\mathrm{l}}$ — the distance a particle diffuses in the dense phase between reactions (analogous to $\ell_{\mathrm{g}}$, Eq. (5)) — the reactions equilibrate at the core of the droplet, so the net loss of sticky material occurs only at its interface and hence the total conversion rate is proportional to droplet surface area. Thus, the dense-phase fluxes are the net reaction rate over droplet area,
$J_{\mathrm{s}\mathrm{l}}(R)=-J_{\mathrm{ns}\mathrm{l}}(R)=\begin{cases}-\frac{1}{3}s_{\mathrm{l}}R,&R\ll\ell_{\mathrm{l}},\\ -s_{\mathrm{l}}\ell_{\mathrm{l}}\left(1-\frac{\ell_{\mathrm{l}}}{\ell_{\mathrm{g}}}\frac{\ell_{\mathrm{g}}}{R}\right),&R\gg\ell_{\mathrm{l}}.\end{cases}$ (7)
By contrast, Ostwald ripening, which opposes this loss of material, occurs with a rate proportional to droplet radius; the gas-phase reactions contribute additional ripening at a rate proportional to the area as well.
With these preparations, we manipulate Eq. (3) into the form
$\frac{\mathrm{d}R}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R}[\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R)-\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)].$ (8)
@@
Eq. (8) is a cruical outcome — upon comparison with Eq. (1), it emphasizes that for arbitrary activity, this system is bound by the scaling laws of conventional Ostwald ripening. This must be so due to the existence of a conserved concentration field; indeed the total concentration $\phi_{\mathrm{s}}+\phi_{\mathrm{ns}}$ is unaffected by activity. Therefore, in our search for acceleration, we learn that the scaling laws expected from Ostwald ripening are the best case scenario, requiring small $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, and our only hope is to increase the rate coefficient. Otherwise, in the “worst case scenario” of large $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, Ostwald ripening can be prevented entirely.
From the second equality of Eq. (3) (the material conservation constraint [28]), we obtain the remaining unknown $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)$ in terms of a set of unitless parameters. Inserting this $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ in Eq. (8), we find the explicit radius growth rate equation akin to Eq. (1), amended to the active case. The values of these unitless parameters determine whether the system achieves acceleration. To obtain acceleration, we require $s_{\mathrm{l}}\ell_{\mathrm{l}}^{2}/D_{\mathrm{g}}\ll(\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}})$ and $s_{\mathrm{l}}\ell_{\mathrm{l}}^{2}/D_{\mathrm{g}}\ll(1+\ell_{\mathrm{g}}/\ell_{\mathrm{l}})(\lambda_{\mathrm{s}\mathrm{g}}/\ell_{\mathrm{g}})$. Otherwise, size control [3] (where a finite, stable droplet radius is reached instead of unlimited growth) or arrested ripening [6] (no shrinkage of small droplets), respectively, occur instead of accelerated ripening; see details in Sec. SII F and Fig. S1. We must further assume that the interplay between dense-phase reactions and thermodynamics do not destabilize the dense phase in the first place [7, 8, 9]; see Sec. SII E.
Supposing both inequalities for the dense-phase activity $s_{\mathrm{l}}\ell_{\mathrm{l}}^{2}/D_{\mathrm{g}}$ hold and the dense phase is stable, we arrive at the active counterpart of Eq. (1),